# Applications and resource estimates for open system simulation on a quantum computer

### v1

@@ III.2 Metal-Insulator transition material search application problem
We suggest starting the search with $t$-$J$ model transport on a 10$\times$10 grid, which we will use in the Landauer setup [landauer1987electrical] (attaching fermion sources and sinks on the left and right edge) as the foundation for the resource estimate. 4$\times$4 grid has been done classically via exact diagonalization and Kubo formula[zemljivc2005resistivity], while 10x10 is large enough to be classically challenging and make the finite size effects negligible. A simulation of transport in an integrable model shows the non-universal effect of sources and sinks to extend only 2-3 sites in [prosen2008third, Fig. 5].
@@
We expect that the material search will start with simple problems like the $10\times 10$ example and eventually go into the ab initio investigation as guided by the experts analyzing the results of the previous runs. Such an investigation may require up to $300$ orbitals per site. We will use the average $150$ orbitals for the resource estimate between the low-end and high-end costs. $300$ orbitals per site were used in [cui2023ab] for ab initio prediction of superconducting order (which then was used to roughly predict $T_{c}$).
Our application instance is a direct generalization of the problem considered [zemljivc2005resistivity]. It is a $t$-$J$ model on a square lattice:
$H=-\sum_{i,j,s}t_{ij}{\tilde{c}}_{js}^{\dagger}{\tilde{c}}_{is}+J\sum_{\langle ij\rangle}({\bf{S}}_{i}{\bf{S}}_{j}-\frac{1}{4}n_{i}n_{j})$ (11)
@@
Each lattice site contains two fermionic modes labeled by $s=\pm 1$. The $\tilde{c}_{i,s}^{\dagger}=(1-n_{i,-s})c_{i,s}^{\dagger}$ are projected creation operators that do not allow the creation of the double-occupancy states, thus satisfying the Hubbard term $Un_{i,s}n_{i,-s}$ perfectly, which is why it is not present in Eq. (11). The perturbative interaction $J=2t^{2}/U$. Here $n_{i,s}=\sum_{s}c^{\dagger}_{i,s}c_{i,s}$ and $S_{i}^{\mu}=\frac{1}{2}\sum_{ss^{\prime}}c^{\dagger}_{i,s}\sigma^{\mu}_{ss^{\prime}}c_{i,s^{\prime}}$, where $\sigma^{\mu}$ are Pauli matrices. The hoppings contain nearest-neighbor and next-nearest neighbor terms $t$ and $t^{\prime}=-0.15t$. The exchange $J=0.3t$ and [zemljivc2005resistivity] take $t=0.3eV$. These choices are typical for downfolded models of cuprates. The $t$-$J$ model is an approximation to the Hubbard model, obtained from the corresponding Hubbard Hamiltonian perturbatively and by restricting to zero and single occupancy subspace of the Hilbert space. In other words, states with two fermions of opposite spin on at least one of the sites are not included in the Hilbert space. The projections are appropriately included in the definition of the creation annihilation operators in the Hamiltonian above. On a quantum computer, we will model the full Hilbert space anyway, so one could instead simulate a simpler Hubbard model. We leave the $t$-$J$ model in this application study for the direct comparison with [zemljivc2005resistivity]. We expect that the finite-size effects will not be as prominent at larger system sizes; therefore, we will use open boundary conditions. The authors compute conductivity via the Kubo formula as a function of frequency $\omega$, averaged over various boundary conditions, and apply broadening by $\epsilon=0.07t$. This broadens the delta functions present due to coherent no-scattering transport. After that broadening, they evaluate the resistivity as $6.6\text{A}/(\text{conductivity}(\omega=0)e_{0}^{2}/\hbar)$ and find good agreement with experimental values.
@@
The initial state of the model above is characterized by a parameter $c_{h}$ describing the number of holes in the system, which is a conserved quantity of the Hamiltonian. For example, $c_{h}=3/20$ means that the system of 20 sites is doped in such a way that three positions are unoccupied, and all the other positions are occupied by 1 electron each. As the initial states are not required to be in equilibrium, any fermionic state with this fraction of excitations can serve as an initial state. For example, one may start with each position single-occupied, and then destroy electrons on the first $c_{h}L^{2}$ sites for an $L\times L$ lattice. To formulate an open system problem equivalent to the Kubo formula used in [zemljivc2005resistivity], we set up sources and sinks for the electrons on two opposite sides of the square lattice, thus breaking the conservation. To derive the Lindblad operators for sources and sinks, one needs to specify the spectral density $\gamma_{L,R}(\omega)$ of the creation-annihilation operators in our system as a function of temperature and chemical potential. This may be done either as a separate task on a quantum computer or classically using a smaller size patch or a free fermion approximation of the system to compute correlation functions and spectral density. Either way, an efficiently computable interpolation for the spectral density needs to be used. One then would use the obtained spectral density to build the Lindblad operators $L_{\mu}^{\text{therm}}$ on the left and right boundary from the hopping terms split by the system boundary, with spectral densities for two different values of chemical potential on the left and the right side, according to the general prescription of [chen2023quantum].
@@
Classical state-of-the-art Recent ab initio numerical investigation by [cui2023ab] shows that even the high-Tc transition temperature can now be predicted by numerical techniques. Since the metal-insulator transition in application-relevant materials happens at higher temperatures, there is a possibility that classical techniques may predict it fully ab initio as well. The apparent absence of publications on this topic may be due to conceptual difficulties with computing a transport coefficient numerically from first principles (unknown disorder, different assumptions of Kubo and Landauer approaches, apparent divergence of DC conductivity without disorder), not due to computational hardness. We also note that machine learning may be successful in the material search for MIT materials, with some initial results presented in [georgescu2021database]. Finally, the cost of searching for MIT transitions experimentally at room temperatures may not be very high for cheap fabrication methods. What makes large throughput experimental material searches unfeasible in condensed matter is that the material properties are highly sensitive to fabrication quality, and such a search would also require automated optimization of fabrication technology for each material. Another concern is that many of the candidate materials are toxic.
@@ IV Quantum Implementation
Many algorithms for tasks related to ours have been proposed. Below we list the promising alternative approaches before describing the approach we believe to be optimal:
  • Thermal Sampling [cubitt2023dissipative] - we found the resource cost estimate of this algorithm hard to apply to the systems we consider.
  • Thermalizing Open Systems with exact fermionic baths [chen2021fast]. While it does not implement every Lindbladian, it covers the ones we need for both of our application instances. We believe that the algorithmic cost of these baths has suboptimal scaling.
@@
What we are interested in is commonly referred to as quantum signal processing (QSP) scaling, which refers to $T$log$(1/\epsilon)$ scaling of the algorithm runtime with evolution time $T$ and precision $\epsilon$, possibly at the cost of extra powers of the system size in front.
IV.1 Lindbladian QSP
We will follow the approach of [chen2023quantum] to achieve QSP scaling for a Lindbladian. First, consider a simple (not achieving QSP scaling) evolution step in the language of [chen2023quantum] via the circuit in Fig. 3.
@@ IV.1 Lindbladian QSP
Here $U$ is a block encoding of a Lindbladian, containing all the Lindblad operators conditioned on the value of the ancilla. Iterating this is the simplest way to simulate the open system dynamics. A more advanced algorithm also found in [chen2023quantum] achieves QSP scaling for the number of gates. Specifically, Theorem III.2 of [chen2023quantum] uses a circuit more advanced than shown above, a compression algorithm on the number of applications of the Lindbladian in the weak measurement version of QSP. To our knowledge, there are no QSP scaling algorithms for time-dependent Lindbladians. Developing them is an important direction for future work. The only QSP scaling algorithm for closed system time-dependent Hamiltonian is [haah2021quantum].
@@ V Resource estimation for proposed quantum algorithms
While Lindbladian QSP may have an optimal scaling (except for the factor of the system size due to block encoding), for smaller system sizes, other algorithms, such as exact fermionic baths, may have an advantage in gate count. As the gate count for both is only available in terms of big-O notation, these algorithms require new resource estimates.
We note that to do the resource estimates for the applications, we need to specify the time to run the algorithm. For the preparation of the equilibrium state, this time is a property of the Lindbladian. In principle, it can be extrapolated from the sizes that can be addressed classically. If those sizes prove too small, one can use the formula in [chen2021fast] relating the relaxation time to the parameters of the eigenstate thermalization hypothesis, which may be determined at smaller sizes.
For Ca3Co2O6 and MIT material applications, we do not actually need to prepare the Gibbs state accurately. The nonequilibrium states seen before the system fully relaxes are the target for quantum simulation of Ca3Co2O6, so the simulation time should match the experiment duration, or at least $1$s of it as the minimum we set. For MIT materials, one approach is to rescale the time corresponding to the broadening 0.07t used in [zemljivc2005resistivity] by a factor of $150\cdot 10/4=375$ due to the increase in the linear size of the system ($L=10$) and the average number of orbitals (150). In other words, $T=375/0.07t$.
We note that [zemljivc2005resistivity] computed the resistivity using the Kubo formula, which would translate into a closed system algorithm that measures the Fourier transform of the correlation function and subsequently integrates it. It may be possible to obtain a comparable performance of this algorithm by performing all the integration inside the quantum algorithm, but naively the open system algorithm requires doing the dynamics just once to obtain the steady state and measure the flow directly, while the closed system algorithm requires performing the dynamics multiple times and integrating the results. We will only do the resource estimate for the open system case.
@@ V.1.1 Lindbladian QSP
Theorem III.2 of [chen2023quantum] lists the following requirements for simulating the open system time evolution for time $t$ to precision $\epsilon$. It relies on the block encoding of the Lindbladian $\mathcal{L}$ governing the evolution, i.e. a unitary $U$ such that all $M$ Lindblad generators $L_{\mu}$ can be found as blocks in the upper left corner of it:
$U=\left(\begin{array}[]{cccc|c}L_{1}&L_{2}&\dots&L_{M}&\cdot\\ \hline\cr\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)$ (12)
Let $c$ be the number of ancillas used by the block encoding of Lindbladian $U$. The quantum algorithm uses $O((c+\text{log}((t+1)/\epsilon))\text{log}((t+1)/\epsilon))$ resettable ancillas, $O((t+1)\text{log}((t+1)/\epsilon)/\text{loglog}((t+1)/\epsilon))$ controlled uses of $U$ and its conjugate, same amount of uses of V that block encodes the Hamiltonian, and $O((t+1)(c+1)\text{polylog}((t+1)/\epsilon))$ other 2-qubit gates.
@@ V.1.2 Block encoding for a Lindbladian
The block encoding used for efficiently implementing $L_{\mu}^{\text{therm}}$ used for Gibbs state preparation is developed in [chen2023quantum]. Those generators rely on the ability to implement the controlled closed system evolution, as well as the custom gates Prepare, $V_{jump}$ that, together with quantum Fourier transform (QFT) allow to integrate that closed system evolution. We do not go into the detail of that construction, providing a high-level overview of the circuit in Fig. 4.
Figure 4: A circuit for the block-encoding of a thermal Lindbladian $U$ used in [chen2023quantum]. The gate $Y_{\delta}=e^{-iY\text{arcsin}\sqrt{\delta}}$ and here $\delta=1-\gamma(\omega)$ is taken conditionally on the frequency of each contribution to the Lindblad operator
The total cost is given in Lemma I.1 of [chen2023quantum]: the maximum closed system controlled simulation time is $T_{CG}$, the internal parameter of $L_{\mu}^{\text{therm}}$, and besides $n$ system qubits extra ancillas logarithmic in the number of jump operators $A$ and in the Fourier discretization grid size $N$ are used. The internal simulation time $T_{CG}$ and grid size $N$ are specified in [chen2023quantum]. We note that for preparing the equilibrium state (in contrast to the open system evolution), this is not the optimal algorithm, and the authors of [chen2023quantum] present further refinements.
For our application instances with simpler Lindbladians, we use the internal structure of $V_{jump}$ in [chen2023quantum]. A set of unitary operators $A^{a}$ is block encoded in it as follows:
$V_{jump}=\left(\sum_{a\in A}|a\rangle\langle a|\otimes A^{a}\right)\cdot(B\otimes I_{\text{sys}}),$ (13)
where the operator $B$ prepares a uniform superposition $B|0\rangle=\sum_{a\in A}\frac{1}{\sqrt{|A|}}|a\rangle$ following [shukla2024efficient]. For the simplest case where $L_{a}=A^{a}$ which are unitary, the block encoding of the Lindbladian is just $U=V_{jump}$ illustrated in Fig. 5.
@@
Since in our case the operators $L_{\mu}$ are not unitary, and have $O(1)$ cases of different coefficients in front, we will need to utilize an $O(1)$ number of extra ancillae for the block encoding.
For Ca3Co2O6, the Lindblad terms (Eq. (6)) contain $\sigma^{\pm}$ multiplied by projectors onto the state of a constant number of neighboring spins. For MIT problem, after the Bravyi-Kitaev transformation [bravyi2002fermionic] from fermions to qubits the creation-annihilation operators would have a logarithmic weight, containing a product of Pauli’s and one $\sigma^{\pm}$. We will now briefly outline a possible (not necessarily optimal) block-encoding, and show that the added complexity is subleading anyways. First we note that $\sigma^{\pm}$ and can be block encoded into a 2-qubit permutation:
$\left(\begin{array}[]{c|c}\sigma^{-}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right)=\left(\begin{array}[]{cc|cc}0&0&1&0\\ 1&0&0&0\\ \hline\cr 0&1&0&0\\ 0&0&0&1\\ \end{array}\right),\quad\left(\begin{array}[]{c|c}\sigma^{+}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right)=\left(\begin{array}[]{cc|cc}0&1&0&0\\ 0&0&1&0\\ \hline\cr 1&0&0&0\\ 0&0&0&1\\ \end{array}\right)$ (14)
A conditioned product $C(U_{1}\otimes U_{2})$ of 2 unitaries $U_{1},U_{2}$ acting on two different sets of qubits can be implemented as a sequence of $CU_{1}$ and $CU_{2}$, which allows to implement the block encoding of creation-annihilation operators.
@@
$\left(\begin{array}[]{ccccc|ccccc}1&0&\dots&0&0&0&0&\dots&0&0\\ 0&0&\dots&0&0&1&0&\dots&0&0\\ 0&0&\dots&0&0&0&1&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&\dots&0&0&0&0&\dots&1&0\\ \hline\cr 0&1&\dots&0&0&0&0&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&\dots&0&1&0&0&\dots&0&0\\ 0&0&\dots&0&0&0&0&\dots&0&1\\ \end{array}\right)$ (15)
Since $k=O(1)$, only a constant number of ancillae is required for the implementation of the MC spin flip rate operators.
Finally, if there are $m=O(1)$ distinct rates in front of $M$ Lindblad terms, we may use additional log $m$ ancillae, prepare a state of it where the amplitudes reflect the desired rates, and then condition $B_{S}$ ($B_{S}|0\rangle=\sum_{a\in S}\frac{1}{\sqrt{|S|}}|a\rangle$) for each subset $S$ of $A$ with the same rate on the state of these ancillae. For MIT material, all $M$ terms have the strength $1/\sqrt{M}$, which effectively requires the increase in the evolution time by a factor of $\alpha=M$. For Ca3Co2O6, the terms that lower the energy have the maximal rate $1$. There is one such term for each spin. We will assume that those terms dominate the amplitude, thus time will effectively increase by a factor $\alpha=2025$, the system size. The total number of terms $M$ is at most $2025\cdot 2^{d}$, where $d$ is the degree of the interaction graph. A more careful estimate of possible neighbors goes as follows. From Fig. 2 we see that there are 2 neighbors linked by $J_{1}$ and 12 neighbors linked by $J_{2}=J_{3}$. There are $m+1$ possible values of the total spin for $m$ spins. Thus the number of possible rates is $3\cdot 13$ (we took into account that half of them lower the energy and thus map to $1$). We will use $M\sim 39\cdot 2025$ the Ca3Co2O6 instance.
The block encoding of $L_{\mu}$ for all $\mu=1\dots M$ only involves the register of $c\sim\text{log}_{2}M+O(1)$ size that is used to control individual $L_{\mu}$. The number of ancillae used by the logical algorithm utilizing this block encoding and Lindbladian QSP is $O((\text{log}M+\text{log}((t+1)/\epsilon))\text{log}((t+1)/\epsilon))$, which is subleading compared to the system size.

### v2

@@ III.2 Metal-Insulator transition material search application problem
We suggest starting the search with Hubbard model transport on a 10$\times$10 grid, which we will use in the Landauer setup [landauer1987electrical] (attaching fermion sources and sinks on the left and right edge) as the foundation for the resource estimate. 4$\times$4 grid has been done classically via exact diagonalization and Kubo formula [zemljivc2005resistivity], while 10x10 is large enough to be classically challenging and make the finite size effects negligible. A simulation of transport in an integrable model shows the non-universal effect of sources and sinks to extend only 2-3 sites in [prosen2008third, Fig. 5].
We note that the Kubo formula computation [zemljivc2005resistivity] can be translated into a closed system quantum algorithm that measures the Fourier transform of the correlation function and subsequently integrates it. This is in contrast to the open system Landauer approach we pursue. It may be possible to obtain a comparable performance of this algorithm by performing all the integration inside the quantum algorithm, but naively the open system algorithm requires doing the dynamics just once to obtain the steady state and measure the flow directly, while the closed system algorithm requires performing the dynamics multiple times and integrating the results. Also, the closed system algorithm will need to start with preparing the thermal state, which may have a comparable cost to our open system evolution. We will only do the resource estimate for the open system case.
@@
We expect that the material search will start with simple problems like the $10\times 10$ example and eventually go into the ab initio investigation as guided by the experts analyzing the results of the previous runs. Such an investigation may require up to $300$ orbitals per site. Indeed, $300$ orbitals per site were used in [cui2023ab] for ab initio prediction of superconducting order (which then was used to effectively predict $T_{c}$). Besides the Hubbard model, we will propose a simple model that can be used as a way to estimate the resources required for the average case of the material search. This model will contain $\sim 150$ orbitals per site as the middle ground between the low-end and high-end costs.
Our Hubbard application instance is a direct generalization of the problem considered in [zemljivc2005resistivity]. While a $t$-$J$ model is considered in that work, it is known to be a large-$U$ approximation of the Hubbard model, which is easier to simulate on a quantum computer. Fermi-Hubbard Hamiltonian describes spin-$1/2$ fermions on a square lattice:
$H=-\sum_{i,j,s}t_{ij}{\tilde{c}}_{js}^{\dagger}{\tilde{c}}_{is}+U\sum_{i}n_{i,\uparrow}n_{i,\downarrow}$ (11)
@@
Each lattice site contains two fermionic modes labeled by $s\in\{\uparrow,\downarrow\}$. The hoppings contain nearest-neighbor and next-nearest neighbor terms $t$ and $t^{\prime}=-0.15t$. The repulsion $U=\frac{20t}{3}$ and [zemljivc2005resistivity] take $t=0.3eV$. These choices are typical for downfolded models of cuprates. We will aim for a direct comparison with [zemljivc2005resistivity]. Though they use the periodic boundary conditions, we expect that the finite-size effects will not be as prominent at larger system sizes; therefore, we will use open boundary conditions. The authors compute conductivity via the Kubo formula as a function of frequency $\omega$ and inverse temperature $\beta$, averaged over various boundary conditions, and apply broadening by $\epsilon=0.07t$. This broadens the delta functions present due to coherent no-scattering transport. After that broadening, they evaluate the resistivity as $6.6\text{A}/(\text{conductivity}(\omega=0)\cdot e_{0}^{2}/\hbar)$ and find good agreement with experimental values.
@@
The initial state of the model above is characterized by a parameter $c_{h}$ describing the number of holes in the system, which is a conserved quantity of the Hamiltonian. For example, $c_{h}=3/20$ means that the system of 20 sites has enough electrons to fill every site with one electron of some spin value, except for three unoccupied positions. As the initial states are not required to be in equilibrium, any fermionic state with this fraction of excitations can serve as an initial state. For example, one may start with each position single-occupied, half in $\downarrow$ and half in $\uparrow$ spin state, and then destroy electrons on the first $c_{h}L^{2}$ sites for an $L\times L$ lattice. To formulate an open system problem equivalent to the Kubo formula used in [zemljivc2005resistivity], we set up sources and sinks for the electrons on two opposite sides of the square lattice, thus breaking the particle-hole symmetry. We note that at half-filling a product of particle-hole and reflection remains a symmetry of the open system dynamics. To derive the thermal Lindblad operators for sources and sinks for general $c_{h}$ and $\beta$, one needs to specify the spectral density $\gamma_{L,R}(\omega)$ of the creation-annihilation operators in our system as a function of temperature and chemical potential. This may be done either as a separate task on a quantum computer or classically using a smaller size patch or a free fermion approximation of the system to compute correlation functions and spectral density. Either way, an efficiently computable interpolation for the spectral density needs to be used. One then would use the obtained spectral density to build the Lindblad operators $L_{\mu}^{\text{therm}}$ on the left and right boundary from the hopping terms split by the system boundary, with spectral densities for two different values of chemical potential on the left and the right side, according to the general prescription of [chen_quantum_2023].
For our resource estimate, we will use the simplest possible case $L_{\text{left}}=c$ and $L_{\text{right}}=c^{\dagger}$ on every site and spin type of the left and right edges of the square lattice, and compare it to the high temperature of the results [zemljivc2005resistivity]. If this system equilibrates for a time $T$, the transport obtained would be roughly equivalent to the broadening of the delta functions in the closed system Kubo formalism used in [zemljivc2005resistivity] by $1/T$. This gives $T=1/\epsilon=1/0.07t$, which we further adjust by a factor $10/4$ of linear increase in the system size:
$T_{\text{min}}\approx 36$ (12)
The full equilibration time may be much longer. The only handle we have on it is the lower bound from the fact that the information propagating linearly should reach the other end of the system, which motivates linear rescaling.
@@
Classical state-of-the-art Recent ab initio numerical investigation by [cui2023ab] shows that even the high-$T_{c}$ transition temperature can now be predicted by numerical techniques. Since the metal-insulator transition in application-relevant materials happens at higher temperatures, there is a possibility that classical techniques may predict it fully ab initio as well. The apparent absence of publications on this topic may be due to conceptual difficulties with computing a transport coefficient numerically from first principles (unknown disorder, different assumptions of Kubo and Landauer approaches, apparent divergence of finite size DC conductivity without disorder), not due to computational hardness. We also note that machine learning may be successful in the material search for MIT materials, with some initial results presented in [georgescu2021database]. Finally, the cost of searching for MIT transitions experimentally at room temperatures may not be very high for cheap fabrication methods. What makes large throughput experimental material searches unfeasible in condensed matter is that the material properties are highly sensitive to fabrication quality, and such a search would also require automated optimization of fabrication technology for each material. Another concern is that many of the candidate materials are toxic.
Average application instance for material search While simulating Eq. (11) will stimulate theory research, it is not directly useful for an ab initio material search. Ab initio prediction [cui2023ab] of superconductivity required keeping 150 orbitals per site without downfolding. For the back-of-the-envelope estimate of the cost of our Landauer approach when used for such ab initio predictions, we assume that a typical material search will involve simulating instances on a 10$\times$10 grid where each site is also described by $\sim 10^{2}$ orbitals. Among ab initio systems of that size, the one with the most optimized quantum simulation algorithms is FeMoCo (one of the latest improvements is in [lee2021even]). We will construct the following artificial Hamiltonian that is meant to have the same features as an average instance of the material search: consider a 10$\times$10 grid of noninteracting FeMoCo Hamiltonians. On each site of the boundaries, Lindblad operators $L_{\mu}^{\text{therm}}$ are applied that prepare a low-temperature state corresponding to an excess or absence of some of the electrons. We have freedom in their choice, and one way to guarantee an approach to appropriate Gibbs distribution is to follow [chen_quantum_2023]. The appropriate norm of the FeMoCo Hamiltonian (not $\|H\|$, but we will use them interchangeably for the back-of-the-envelope estimate) is $\lambda\sim 1500$ [lee2021even], and the cost of its block-encoding is $\sim 17,000$ [harrigan2024expressing]. We rescale the required evolution time by $150$ orbitals:
$T_{\text{Av.}}\sim 5000$ (13)
The estimates for this instance will quickly result in very high resource costs, so we leave it for future work to develop new simpler forms of $L_{\mu}^{\text{therm}}$ that have low-temperature Gibbs state as the steady state and algorithms compatible with this model.
@@ IV Quantum Implementation
For both applications we consider, we require an algorithm that would simulate an evolution (for time $T$, to precision $\epsilon$) of state $\rho$ under an open system dynamics given by a Hamiltonian $H$ and Lindblad generators $L_{i}$. In both cases, $H$ is a sum of local terms, and $L_{i}$ are local. Many algorithms for this task have been proposed. We will briefly mention other approaches before describing the approach we believe to be optimal:
@@
  • Any approximate implementation of $e^{\delta\mathcal{L}_{i}}$ for a small $\delta$ and the Lindbladian with one generator $L_{i}$ can be used in a Trotter formula, which we do in App. F for the example of an exact resource count calculation. The scaling with time is $\sim T^{2}$. The power is improved in [ding2024single], approaching $1$ as the order of the Trotter formula is increased.
What we are interested in is commonly referred to as quantum signal processing (QSP) scaling, which is the $T$log$(1/\epsilon)$ scaling of the algorithm runtime with evolution time $T$ and precision $\epsilon$, possibly at the cost of extra powers of the system size in front.
Lindbladian QSP: QSP scaling for open system dynamics was originally achieved in [cleve2016efficient]. We will follow the approach of [chen_quantum_2023] that allows further optimizations during the block encoding of $\mathcal{L}$.
First, consider a simple (not achieving QSP scaling) evolution step in the language of [chen_quantum_2023] via the circuit in Fig. 3.
@@
Here $U$ is a block encoding of a Lindbladian containing all the Lindblad operators conditioned on the value of the ancilla. Iterating this is the simplest way to simulate the open system dynamics. A more advanced algorithm also found in [chen_quantum_2023] achieves QSP scaling for the number of gates. Specifically, Theorem III.2 of [chen_quantum_2023] uses a circuit more advanced than shown above, a compression algorithm on the number of applications of the Lindbladian in the weak measurement version of QSP. To our knowledge, there are no QSP scaling algorithms for time-dependent Lindbladians. Developing them is an important direction for future work. A QSP scaling algorithm for a closed system time-dependent Hamiltonian on a lattice was demonstrated in [haah2021quantum].
While Lindbladian QSP may have an optimal scaling (except for the factor of the system size due to block encoding), for small errors and evolution times other algorithms, such as Trotter, may have an advantage in gate count. As the gate count for both is only available in terms of big-O notation, these algorithms require new exact resource estimates. Our main results rely on order-of-magnitude estimates that do not keep track of the exact prefactor. We also present preliminary, heavily suboptimal exact resource estimates for Trotter in App. F.
@@ V.1.1 Lindbladian QSP
Theorem III.2 of [chen_quantum_2023] lists the following requirements for simulating the open system time evolution for time $t$ to precision $\epsilon$. It relies on the block encoding of the Lindbladian $\mathcal{L}$ governing the evolution, i.e., a unitary $U$ such that all $M$ Lindblad generators $L_{\mu}$ can be found as blocks in the upper left corner of it, and the block encoding $V$ of the system Hamiltonian:
$U=\left(\begin{array}[]{c|c}L_{1}&\cdot\\ L_{2}&\cdot\\ \dots&\cdot\\ L_{M}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right),\quad V=\left(\begin{array}[]{c|c}H&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right)$ (14)
Let $c$ be the number of ancillas used by $U$ and $V$. The quantum algorithm uses $O((c+\text{log}((t+1)/\epsilon))\text{log}((t+1)/\epsilon))$ resettable ancillas, $O((t+1)\text{log}((t+1)/\epsilon)/\text{loglog}((t+1)/\epsilon))$ controlled uses of $U$ and its conjugate, same amount of uses of each of $V$ that block encodes the Hamiltonian, and $O((t+1)(c+1)\text{polylog}((t+1)/\epsilon))$ other 2-qubit gates. It remains unspecified how many of those gates are Clifford, as the details of the composition of the circuit outside $U$ are not fully presented in [chen_quantum_2023], instead appearing in various references therein. We note that there can be no more than $O((t+1)(c+1)\text{polylog}((t+1)/\epsilon))$ non-Clifford gates and arbitrary rotations outside $U$. As we will be able to perform a detailed estimate of the cost of (controlled) $U$, it will become apparent that the non-Clifford count $O((t+1)(c+1)\text{polylog}((t+1)/\epsilon))$ is subleading. Our strategy will be to keep only the leading terms in the cost and replace big-O notation with a factor 1. This will be sufficient for the initial analysis we perform. As an example of rigorous resource estimation, we analyze a suboptimal evolution step in Fig. 3 in a Trotter evolution algorithm in App. F.
@@ V.1.2 Block encoding for a Lindbladian
The general way to construct $U$ defined in Eq. (14) follows the structure of what is denoted $V_{jump}$ in [chen_quantum_2023]. A set of unitary operators $A^{a}$ is block encoded in it as follows:
$V_{jump}=\left(\sum_{a\in A}|a\rangle\langle a|\otimes A^{a}\right)\cdot(B\otimes{\mathds{1}}_{\text{sys}}),$ (15)
where the operator $B$ prepares a uniform superposition $B|0\rangle=\sum_{a\in A}\frac{1}{\sqrt{|A|}}|a\rangle$ following [shukla2024efficient]. For the simplest case where $L_{a}=A^{a}$, which are unitary, the block encoding of the Lindbladian is just $U=V_{jump}$ illustrated in Fig. 4.
@@
Since in our case the operators $L_{\mu}$ are not unitary and have $O(1)$ cases of different coefficients in front, we will need to utilize an $O(1)$ number of extra ancillae for the block encoding:
$U=\left(\sum_{\mu=1}^{M}|\mu\rangle\langle\mu|\otimes U_{\mu}\right)\cdot(B\otimes{\mathds{1}}_{\text{anc}}\otimes{\mathds{1}}_{\text{sys}})$ (16)
where $U_{\mu}$ is a block encoding of an individual Lindblad operator $L_{\mu}$:
$U_{\mu}=\left(\begin{array}[]{c|c}L_{\mu}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right).$ (17)
@@
The cost that will dominate is the leading term of system-size dependence coming from the operator $\sum_{\mu=1}^{M}|\mu\rangle\langle\mu|\otimes U_{\mu}$, where $U^{a}$ is the block-encoding of the Lindblad generator $L_{a}$. Let us briefly describe the other operator $B$ and the related rescaling of the Lindblad generators. Depending on the norms of the operators $\|H\|,\|L_{\mu}\|$, rescaled version $H/\alpha_{H},L_{\mu}/\alpha_{L}$ appear in the block encoding Eq. (14). The time has to be rescaled as $t\to\alpha t,~{}\alpha=$max$(\alpha_{H},\alpha_{L})$. We will now compute the rescaling factors $\alpha_{H},\alpha_{L}$. The Hamiltonian one is determined from unitarity of $V$, thus $|\langle 0|V|0\rangle|\leq 1$ and we can take $\|H\|/\alpha_{H}=1$. Another comes from the specific construction above. Let $U_{\mu}$ encode $L_{\mu}/\|L_{\mu}\|$. The operator $B$ contains a matrix element $\sum_{\mu}c_{\mu}|\mu\rangle\langle 0|$ such that $\sum_{m}|c_{\mu}|^{2}=1$. The $U|0\rangle=\sum_{\mu}c_{\mu}/\|L_{\mu}\||\mu\rangle\otimes U_{\mu}$ has the following block structure:
$U=\left(\begin{array}[]{c|c}\frac{c_{1}}{\text{max}(1,\|L_{1}\|)}L_{1}&\cdot\\ \frac{c_{2}}{\text{max}(1,\|L_{2}\|)}L_{2}&\cdot\\ \dots&\cdot\\ \frac{c_{M}}{\text{max}(1,\|L_{M}\|)}L_{M}&\cdot\\ \hline\cr\cdot&\cdot\\ \end{array}\right)$ (18)
The solution is to pick $c_{\mu}=\frac{1}{\alpha_{L}}\|L_{\mu}\|$, which leads to
$\alpha_{L}^{-2}\sum_{\mu}\|L_{\mu}\|^{2}=1$ (19)
The time rescaling factor is then:
$t\to\alpha t,\quad\alpha=\text{max}(\|H\|,\sqrt{\sum_{\mu}\|L_{\mu}\|^{2}})$ (20)
The Hamiltonian contribution is larger for both of our application instances. Ca3Co2O6, $\|H\|\sim 3.2\cdot 2025$, $\alpha\approx 6500$ For Hubbard, $\|H\|\sim 11\cdot 100$, $\alpha\approx 1100$.